(a) You roll a normal dice and score 7. (b) You fall off your bike on the way home from school. (c) You complete all your maths homework correctly.
probability solutions
30 verified PDFs
What is the probability of getting at most one tail? 6) How many even 4 digit numbers are there? 7) Jane and John are preparing for the SHSAT. Jane gets math ...
Hence the probability of at least one match is 1 − 0.0173 = 0.9827. 13. QUESTION: An examination consists of multiple-choice questions, each having five ...
What is the probability that the last person to board gets to sit in his assigned seat? 2. [D. Knuth] Mr. Smith works on the 13th floor of a 15 floor building.
Probability of an event E = Number of outcomes favourable to E. Total number of outcomes. 3. 0 ≤ P (E) ≤ 1. 4. If P(E) = 0 then it is an impossible event. 5.
What to compute: PA: probability A wins; PB: probability B wins;. PAB: probability of a draw; Q: probability the duel never ends. Page 6. The Solution. Sergey ...
Find the smallest positive integer n such that the probability that the two selected squares are horizontally or vertically adjacent is less than 2015. 1. Remark: This problem was from AIME 2015. 3. Five red and five green cards are shuffled and top five cards are laid out in a row.
Call Alice's number A and Bob's number B. The is the probability that A ≤ B? What is the probability that A + B ≤ 7? What is the probability that A − B ...
We toss a fair coin twice. What is the probability that: (a) We get heads then tails. (b) We get two tails. (c) The first toss turns up heads.
There is a 60% chance of rain today. If it rains today, there is a 70% chance it will rain tomorrow. What is the probability that it will rain today and.
